the empirical formula for a compound comprised of 52.14% C, 13.13% H, and 34.73% O contains carbon atoms, hydrogen atoms, and oxygen atoms


1
Expert's answer
2021-01-25T04:22:56-0500

No. of moles of C = 52.14/12 = 4.345

No. of moles of H = 13.13/1 = 13.13

No. of moles of O = 34.73/16 = 2.170

Ratio of C:H:O = 4.345:13.13:2.170 = 2:6:1

So , empirical formula for a compound

= C2H6O